# Study of dark counts in optical superconducting transition-edge sensors

### v1

@@ Abstract
Contrary to initial expectations, the photon-like DCR in the eV energy range remained constant (within uncertainties) across all runs (see the blue points in Fig. 6), despite the introduction of radioactive sources in Run 2 and Run 3. This consistency in the photon-like DCR, regardless of the presence of radioactive sources, suggests that high-energy gammas do not contribute to the observed photon-like event rate. Despite including the Optical Physics package in our GEANT4 simulation, no optical photons indicative of scintillation processes around the TESs were recorded.
@@
One potential explanation for the observed photon-like DCR is the presence of stray optical photons within the cryostat, possibly from external ambient light. However, the likelihood of these photons entering through the optical fiber fixed on the -stage seems minimal, as any ambient light entering the fiber would be absorbed by the -stage. Supporting this, previous research indicates that the presence or absence of a fiber fed into the cryostat, but not directly coupled to the TES, has no significant effect on the photon-like DCR (see Fig. 4.13 in Ref. [14]). This suggests that the source of these photons remains unknown and requires further investigation.
@@
In Run 1, we also selected high-energy signals to calculate the high-energy dark count rate. This rate is considered to be caused by cosmic ray interactions with the TES substrate. Notably, for TES B, whose substrate area is half the size of TES E’s, we observed that the high-energy DCR was half that of TES E, substantiating the relationship between the substrate’s size and the high-energy DCR (see appendix for more details). The high-energy event rate from the cosmic ray simulation and the experimental data do not align within uncertainty. However, the rates are within the same order of magnitude. The observed discrepancies can be attributed to limitations in the CRY package, the simulation framework, and the modeling of our experimental setup.
@@
Upon introducing radioactive sources, a noticeable increase in the rate of high-energy signals in the TES was observed on the oscilloscope. This was particularly visible with the Na source. Our simulation shows that we can accurately model the high-energy events in the presence of a radioactive source, with compatible results for both the Th and Na run. The simulation indicates that the substrate absorbs the energy and then transmits it to the TES, where it is detected. Notably, altering the substrate area results in corresponding changes in the DCR, further confirming this interaction mechanism.
@@
Run 4 provided confirmation that cosmic ray showers are capable of producing high-energy events in the TES, as evidenced by the quadruple coincidences observed between the sensor and the three sabers below the ADR (see Fig. 7). The simulated ratio of triple coincidences to quadruple coincidences overestimates that observed in Run 4A (2337 in simulation and 477 in real life), whereas the simulated ratio of high-energy TES hits to quadruple coincidences underestimates Run 4B (198 in simulation and 321 in real life). This is likely due to the low statistics of quadruple coincidences in both runs.
@@
Conclusions. We have characterized the dark count rate of a Ti/Au superconducting transition-edge sensor. By employing singular-value decomposition, principal component analysis, and K-means clustering, we effectively filtered out noise and high-energy events to isolate photon-like signals. These photon-like signals are of particular interest since they resemble the expected dark photon signal in a dielectric haloscope. We achieved a photon-like DCR of Hz in the eV range. Removing optical connections between the TES and ambient is essential to achieve very low DCRs. The source of the observed residual photon-like events still remains unexplained, necessitating further investigation.
In relation to high-energy events, our runs using Th and Na sources, coupled with a comprehensive GEANT4 simulation, have allowed us to ascribe these events to high-energy depositions in the TES substrate. This marks the first time such a clear attribution has been possible. Additionally, natural radiation appears to have a negligible impact on the TES’s high-energy DCR.
@@
Acknowledgments. We thank Stefano Nisi and Francesco Ferella at Laboratori Nazionali del Gran Sasso (LNGS) for the ICP-MS measurements. This research was carried out on the High Performance Computing resources at New York University Abu Dhabi. We are grateful to the NYUAD Kawader Research Assistantship Program. This work was also made possible by the contribution of Grant (62313) from the John Templeton Foundation. Fermilab is operated by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the United States Department of Energy.
@@ Appendix A TES fabrication
Both TES B and TES E were fabricated at INRiM. They are composed of a film of Ti under a film of Au deposited on a substrate [15]. The substrate was made through low pressure chemical vapor deposition (LPCVC) and is composed of a Si substrate (525 thick), sandwiched between two bilayers composed of SiN silicon nitride ( thick) and SiO thermal oxide ( thick), with the silicon nitride as the outside layer.
@@
Before the deposition, the substrate underwent sputter etching for 15 seconds at . Then, the sample was moved to the UHV chamber for the Ti and Au deposition. This consisted of Ti deposition by e-beam at a rate of followed, after about , by the deposition of Au by effusion cell at a rate of . The deposited thicknesses were continually monitored by a quartz crystal sensor.
@@ Appendix B Classification algorithm
To classify the event type of each waveform, we first apply a Butterworth-filter and calculate a set of variables including maximum amplitude, area, full-width at half maximum, standard deviation, minimum and maximum gradient, and the cross correlation with a reference photon-like event. This gives us a point in an n-dimensional space for each TES pulse recorded by the oscilloscope.
We then perform a principal component analysis on the matrix containing the aforementioned n-dimensional points [10]. Using this technique, we can identify the two axes in this n-dimensional space that maximize variance among the points. Plotting pulses on these two axes typically results in distinct clusters. This representation enables the visual differentiation of photon-like, high-energy, and electrical noise events.
Furthermore, applying K-means clustering allows us to easily collate pulses that fall close together in the n-dimensional space [11]. The combined use of these tools, and, if necessary, reanalysis of selected pulse subgroups, enables efficient pulse shape discrimination on datasets comprised of thousands to tens of thousands of pulses in approximately ten minutes.
@@ Appendix C GEANT4
All code used for the simulations is available on our GitHub repository [18]. We used the GEANT4 v.11.1.2 software, an object-oriented framework written in C++ that employs Monte Carlo methods [19]. In our simulations, we activated the EmLiverMore low-energy physics module with the Particle Induced X-ray Emission (PIXE) parameter enabled.
During each simulation event, data was collected on particle characteristics, including position, initial energy, primary vertex, the sensitive volume or volumes that it hits, and the energy deposited. To simulate cosmic ray particles, which include neutrons, protons, gammas, electrons, pions, and muons, we used the Cosmic Ray Shower Library (CRY) [13].
For simulations involving thorium and sodium, we employed the Radioactive Decay module along with the General Particle Source. To optimize computing time, we limited our analysis to secondary products with a Track ID of 50 or less. We validated this approach by comparing the Th spectrum from our simulation with measurements from the High Purity Germanium detector at NYUAD (Fig. 8).
The geometry of the detector was defined using GDML files, which, when combined with macro files, allowed for the dynamic adjustment of various simulation parameters. For computationally intensive simulations, we used the Message Passing Interface (MPI) on a high-performance computer, which allowed us to use approximately 8000 CPU hours per simulation.
@@ Appendix D TES B
Figure 9 shows the energy calibration curve for TES B.
A background run of 14 hours and 45 minutes resulted in 7 photon-like counts and 370 high-energy counts. The high-energy dark count rate is half that of TES E, which is consistent with the hypothesis that high-energy events are caused by cosmic ray particle interactions within the substrate. Given that the substrate size of TES B is half that of TES E, this reduction in dark count rate is in line with expectations. However, the photon-like DCR is effectively the same as that of TES E, implying that the photon-like events do not originate in the substrate and as such are unaffected by the difference in the size of their substrates.
The fact that the results from both TESs are consistent indicates that the fabrication techniques we employ are reliable and lead to reproducible sensors.
@@ Appendix E Sensitive volume
To assess whether the TESs were sensitive just to energy depositions in the substrate or also to energy depositions in the base of the copper box hosting them, a thorium source experiment involving both TESs was conducted. A GEANT4 simulation was also performed.
@@
Table 1 summarizes our results. Both the experimental data and the simulation show comparable ratios between TES E and TES B counts, as well as similar numbers of double coincidences. These results reinforce our belief that the substrate, which is the sensitive volume in the simulation, is the only volume where energy depositions lead to detections in the TESs.
Given the base of the copper box is not a sensitive volume in the simulation, if energy deposited in the copper could be detected by the TESs in real life, we would expect the simulation to underestimate TES hits. Furthermore, since the base of the copper box connects both TESs, if the copper were able to diffuse heat to the sensors, we would observe significantly more double coincidences in real life than in simulation. Since we did not observe either, we conclude that energy deposited in the copper does not contribute to TES hits.
To investigate further the nature of the double coincidences, we calculated the rate of random coincidences using the formula , where is the number of double coincidences, and are the numbers of hits on TES E and TES B, respectively, T is the gate time, and t is the total measurement time. For our specific setup, which included a gate time of and a total time of 17.5 hours, the expected number of random double coincidences is 0.001. This number is markedly lower than the 7 occurrences we observed over the 17.5 hours. This confirms that the double coincidences are correlated and likely caused by the same initial Th decay. Through simulation, it was understood that the main source of such coincidences is secondary electrons originating in the copper box from Th decay events, impacting the substrates simultaneously.
To further validate that the energy depositions causing the high-energy events occur only within the substrate, we conducted a series of simulations varying both the size of the substrate and its distance from the radioactive source. These simulations conform best to reality when the sensitive volume matches the real substrate size and revealed a square law relationship between the count rate and the distance from the radioactive source.
@@ Appendix F Saber calibration
The calibration of the sabers was carried out at NYUAD using different radioactive sources. Fig. 11 shows the calibration curve for all three sabers with the PMTs operating at similar gains.

### v4

@@ VI Results and discussion
Contrary to initial expectations, the photonlike DCR in the 0.8–3.2 eV energy range remained constant (within uncertainties) across all runs (see the blue points in Fig. 6), despite the introduction of radioactive sources in Runs 2 and 3. This consistency in the photonlike DCR, regardless of the presence of radioactive sources, suggests that gamma rays do not contribute to the observed photonlike event rate. Despite including the Optical Physics package in our geant4 simulation, no optical photons indicative of scintillation processes around the TESs were recorded.
@@
One potential explanation for the observed photonlike DCR is the presence of stray optical photons within the cryostat, possibly from external ambient light. However, the likelihood of these photons entering through the optical fiber seems minimal, as any ambient light entering the fiber would be absorbed by the 3-K stage to which it is fixed. Supporting this, previous research indicates that the presence or absence of a fiber fed into the cryostat, but not directly coupled to the TES, has no significant effect on the photonlike DCR (see Fig. 4.13 in Ref. [20]). This suggests that the source of these photonlike signals remains unknown, hence the absence of simulation points in Fig. 6 for the photonlike DCR, and requires further investigation.
@@
In Run 1, we selected high-energy signals to calculate the high-energy dark-count rate. This rate is caused by cosmic ray and environmental gamma-ray interactions with the TES substrate, as confirmed by our simulations. Notably, for TES B, the substrte area of which is half the size of that of TES E, we observed that the high-energy DCR was half that of TES E, substantiating the relationship between the size of the substrate and the high-energy DCR (for more details, see Appendix E).
@@
Upon introducing radioactive sources, a noticeable increase in the rate of high-energy signals in the TES was observed on the oscilloscope. This was particularly visible with the 22Na source. Our simulation shows that we can accurately model the high-energy events in the presence of a radioactive source, with compatible results for both the 232Th and 22Na run. The simulation indicates that the substrate absorbs the energy and then transmits it to the TES, where it is detected. Notably, altering the substrate area linearly changes the DCR, further confirming this interaction mechanism.
@@
Run 4 provided confirmation that cosmic ray showers are capable of producing high-energy events in the TES, as evidenced by the quadruple coincidences observed between the sensor and the three sabers below the ADR (see Fig. 7). The simulated ratio of triple coincidences to quadruple coincidences, which represents the fraction of cosmic ray events detected by the saber system that also deposited energy in the TES, overestimates the one observed in Run 4A by a factor of 5. Similarly, the simulated ratio of high-energy TES hits to quadruple coincidences overestimates Run 4B by a factor of 1.7. These discrepancies are likely due to systematics in the cosmic ray simulation. Table 1 shows the raw counts for triple and quadruple coincidences in simulation and real life.
| Sim | Exp | Ratio Sim/Exp | |
| Run 4A t/q | /46 | 3818/8 | 5 |
| Run 4B h/q | /46 | 963/3 | 1.7 |
VII Conclusions
@@
We have characterized the dark-count rate of a Ti/Au superconducting transition-edge sensor. By employing singular-value decomposition, principal component analysis, and k-means clustering, we have effectively filtered out noise and high-energy events to isolate photonlike signals. These photonlike signals are of particular interest since they resemble the expected dark photon signal in a dielectric haloscope. We have achieved a photonlike DCR of Hz in the eV range. Removing optical connections between the TES and ambient is essential to achieve very low DCRs. The source of the observed residual photonlike events still remains unexplained, necessitating further investigation.
Regarding high-energy events, our runs using 232Th and 22Na sources, along with coincidence measurements between the TES and an external cosmic ray detector, coupled with a detailed geant4 simulation, have allowed us to ascribe these events to high-energy particle impacts in the TES substrate. Previous studies hypothesized that these high-energy events are due to radiation and cosmic ray interactions with the substrate, but have lacked experimental and simulation verification. Additionally, natural radiation from materials in close proximity to the TES appears to have a negligible impact on the high-energy DCR of the TES. Instead, the primary contributors to this rate are cosmic ray particles and environmental gamma rays hitting the substrate. All code used for the simulations is available on our GitHub repository [21].
Acknowledgments
@@
Part of this work has been carried out at QR Lab-Micro&Nanolaboratories, Istituto Nazionale di Ricerca Metrologica (INRiM). We thank Stefano Nisi and Francesco Ferella at Laboratori Nazionali del Gran Sasso (LNGS) for the ICP-MS measurements. This research was carried out on the High Performance Computing resources at New York University Abu Dhabi. We are grateful to the NYUAD Kawader Research Assistantship Program. This work was also made possible by the contribution of Grant (62313) from the John Templeton Foundation. Fermilab is operated by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the United States Department of Energy.
Appendix
@@ Appendix A TES fabrication
Both TES B and TES E were fabricated at INRiM. They are composed of a 15-nm film of Ti under a 30-nm film of Au deposited on a substrate [22]. The substrate was made through low-pressure chemical vapor deposition (LPCVC) and is composed of a Si substrate (525 thick), sandwiched between two bilayers composed of SiNx silicon nitride ( thick) and SiO2 thermal oxide ( thick), with the silicon nitride as the outside layer.
@@
Before the deposition, the substrate underwent sputter etching for 15 s at . Then, the sample was moved to the UHV chamber for the Ti and Au deposition. This consisted of Ti deposition by beam at a rate of , followed, after about , by the deposition of Au by effusion cell at a rate of . The deposited thicknesses were continually monitored by a quartz-crystal sensor.
@@ Appendix B Classification algorithm
To classify the event type of each waveform, we first apply the SciPy Butterworth filter with low-pass mode, order 2, frequency 1/30, which, when applied to the TES triggered data sampled at , corresponds to , and we then calculate a set of variables including the maximum amplitude, the area, the full-width at half maximum, the standard deviation, the average positive and negative gradient, and the cross-correlation with a reference photonlike event. This gives us a point in an n-dimensional space for each TES pulse recorded by the oscilloscope.
We then perform a PCA on the matrix containing the aforementioned n-dimensional points [15]. Using this technique, we can identify the two axes in this n-dimensional space that maximize the variance among these points. Plotting pulses on these two axes typically results in distinct clusters. This representation enables the visual differentiation of photonlike, high-energy, and electrical-noise events.
Furthermore, applying k-means clustering allows us to easily collate pulses that fall close together in the n-dimensional space [16]. The combined use of these tools, and, if necessary, reanalysis of selected pulse subgroups, enables efficient pulse-shape discrimination on datasets comprised of thousands to tens of thousands of pulses in approximately 10 min.
@@ Appendix C GEANT4
We used the geant4 v.11.1.2 software, an object-oriented framework written in C++ that employs Monte Carlo methods [24]. In our simulations, we activated the EmLiverMore low-energy physics module, with the particle-induced x-ray Emission (PIXE) parameter enabled.
During each simulation event, data were collected on particle characteristics, including the position, the initial energy, the primary vertex, the sensitive volume or volumes that it hits, and the energy deposited. To simulate cosmic ray particles, which include neutrons, protons, gamma rays, electrons, pions, and muons, we used the Cosmic Ray Shower Library (CRY) [18]. To simulate the environmental radioactivity we used the gamma spectrum measured in the laboratory, by means of a NaI detector. The gamma rays were generated isotropically on a cylindrical surface surrounding the ADR.
For simulations involving thorium and sodium, we employed the Radioactive Decay module along with the General Particle Source. To optimize the computing time, we limited our analysis to secondary products with a Track ID of 50 or less. We validated this approach by comparing the 232Th spectrum from our simulation with measurements from the High Purity Germanium detector at NYUAD (Fig. 8).
The geometry of the detector was defined using Geometry Description Markup Language (GDML) files, which, when combined with macro files, allowed for the dynamic adjustment of various simulation parameters. For computationally intensive simulations, we used the message passing interface (MPI) on a high-performance computer, which allowed us to use approximately 8000 CPU hours per simulation.
@@ Appendix D TES B
In Fig. 9, we show the energy-calibration curve for TES B.
A background run of 14 hours and 45 minutes resulted in seven photonlike counts and 370 high-energy counts. The high-energy dark-count rate is half that of TES E, the substrate of which is twice the size of that of TES B. This is consistent with the hypothesis that high-energy events are caused by cosmic ray particle interactions within the substrate. However, the photonlike DCR is effectively the same as that of TES E, implying that the photonlike events do not originate in the substrate and, as such, are unaffected by the difference in substrate size.
The fact that the results from both TESs are consistent indicates that the fabrication techniques that we employ are reliable and lead to reproducible sensors.
@@ Appendix E Sensitive volume
To assess whether the TESs were sensitive just to energy depositions in the substrate or also to energy depositions in the base of the copper box hosting them, a thorium source experiment involving both TESs was conducted. A geant4 simulation was also performed (see Fig. 10).
@@
Our results are summarized in Table 2. Both the experimental data and the simulation show comparable ratios between the TES E and TES B counts, as well as similar numbers of double coincidences. These results reinforce our belief that the substrate, which is the sensitive volume in the simulation, is the only volume where energy depositions lead to detections in the TESs.
Given that the base of the copper box is not a sensitive volume in the simulation, if energy deposited in the copper could be detected by the TESs in real life, we would expect the simulation to underestimate the TES hits. Furthermore, since the base of the copper box connects both TESs, if the copper were able to diffuse heat to the sensors, we would observe significantly more double coincidences in real life than in simulation. Since we did not observe either, we conclude that energy deposited in the copper does not contribute to TES hits.
To investigate further the nature of the double coincidences, we have calculated the rate of random coincidences using the formula , where is the number of double coincidences, and are the numbers of hits on TES E and TES B, respectively, T is the gate time, and t is the total measurement time. For our specific setup, which included a gate time of and a total time of 17.5 hours, the expected number of random double coincidences is 0.001. This number is markedly lower than the seven occurrences that we observed over the 17.5 hours. This confirms that the double coincidences are correlated and likely caused by the same initial 232Th decay. Through simulation, it was understood that the main source of such coincidences is secondary electrons, originating in the copper box from 232Th decay events, impacting the substrates simultaneously.
To further validate that the energy depositions causing the high-energy events occur only within the substrate, we have conducted a series of simulations varying both the size of the substrate and its distance from the radioactive source. These simulations conform best to reality when the sensitive volume matches the real substrate size and position and they reveal a square-law relationship between the count rate and the distance from the radioactive source.
@@ Appendix F Saber calibration
The calibration of the sabers was carried out at NYUAD using different radioactive sources. For each source, we identified the characteristic Compton edge and measured the corresponding peak height in the detected amplitude spectrum (in millivolts). In Fig. 11, we show the calibration curve for all three sabers with the PMTs operating at similar gains, while in Table 3, we list the raw data.
| Saber 1 [mV] | Saber 2 [mV] | Saber 3 [mV] | Compton edge [keV] | |
| 133Ba | 78 | 115 | 116 | 207 |
| 137Cs | 240 | 310 | 270 | 478 |
| 22Na | 550 | 620 | 604 | 1062 |
| 232Th | 980 | 945 | 935 | 2381 |
| Muons | 1400 | 1300 | 1210 | 3760 |
Appendix G Simulation and experimental data
In Table 4, we report the values obtained in the simulation and in real life for the photonlike and the high-energy DCR during the background, thorium, and sodium runs.
| Photonlike [mHz] | High-energy [mHz] | High-energy Sim [mHz] | |
| Background | 0.36 | 12.10 | 12.83 |
| 232Th | 0.52 | 34.00 | 39.20 |
| 22Na | 0.00 | 131.00 | 129.00 |